# Perihelion Asymmetry in the Water Production Rate of the Interstellar Object 3I/ATLAS

### v1

@@ Abstract
3I/ATLAS is an interstellar object whose activity provides critical insights into its composition and origin. However, due to its orbital geometry, the object is too close to the Sun near perihelion to be observed from the ground, and space-based measurements are therefore required. Here we characterize the water production rate of 3I/ATLAS using SOHO/SWAN Lyman- observations from 2025 November to December (heliocentric distances 1.4 to 2.2 au) with 3D Monte Carlo modeling. We report a peak post-perihelion water production rate of mol s-1, corresponding to a minimum active fraction of 30% (assuming a maximum nucleus radius of 2.8 km). Comparison of our post-perihelion measurements with published pre-perihelion results reveals a heliocentric asymmetry, with an scaling for the inbound rise, followed by a shallower scaling during the outbound decline, where is heliocentric distance. The post-perihelion behavior indicates that the water production of 3I/ATLAS was driven primarily by the varying solar insolation acting on a stable active area. Combined with other evidence, including comparison with the hyperactive comet 103P/Hartley 2, our findings suggest that its water production is likely dominated by a distributed source of icy grains. Furthermore, it displayed remarkable stability in the activity with no signs of outbursts or rapid depletion of water production.
@@ II.2 Monte Carlo Particle Trajectory Model (MCPTM)
The water production rate was determined by scaling the observed Lyman- integrated flux to the MCPTM predictions. Following the methodology of Combi and Smyth (1988a), the relationship is expressed as:
where is the integrated flux within the photometric aperture, is the SOHO-comet distance, and is the calibration factor (Section II). The term represents the modeled number of hydrogen atoms within the aperture, calculated based on the adopted velocity field () and lifetimes ().
@@ III.1 Post-perihelion phase
We derived water production rates from SOHO/SWAN observations using the MCPTM described in Section II.2. During the post-perihelion phase, the water production rate reached a peak of mol s-1 (Table 1) and declined steadily with a heliocentric index of , as shown in Figure 3.
@@ III.2 Perihelion Asymmetry
Comparison between our post-perihelion measurements and the published pre-perihelion data reveals a significant perihelion asymmetry in water production. The activity profile is characterized by a steep pre-perihelion rise followed by a shallower post-perihelion decline (Figure 3). This asymmetry is also evident in the heliocentric light curve obtained from the COBS database (COBS International Database, 2018).
@@ Appendix A Validation of the Water Production Rate Retrieval Methodology
To validate our data reduction and modeling pipeline, we benchmarked our results against three well-characterized comets: 41P/Tuttle-Giacobini-Kresák (2017 apparition), 46P/Wirtanen (2018 apparition), and C/2020 F3 (NEOWISE). These targets span nearly three orders of magnitude in water production rate ( to mol s-1), allowing for a rigorous assessment of our analysis across a wide dynamic range. As illustrated in Figures 5, 6, and 7, our derived water production rates show excellent agreement with independent space- and ground-based measurements, confirming the robustness of our methodology.
41P/Tuttle-Giacobini-Kresák: As a typical Jupiter-family comet, 41P serves as a benchmark for the sensitivity of our analysis in the low-activity regime ( mol s-1). As shown in Figure 5, our derived rates show excellent agreement with both the space-based Swift/UVOT results (Bodewits et al., 2018) and ground-based TRAPPIST photometry (Moulane et al., 2018). This concordance between the SWAN Lyman- (atomic hydrogen) and TRAPPIST OH data validates the photochemical scale lengths and calibration parameters adopted for low-production targets. 46P/Wirtanen:
@@
C/2020 F3 (NEOWISE): For the highly active Oort Cloud comet C/2020 F3 (NEOWISE), our derived water production rates closely track the heliocentric evolution reported in the standard SWAN analysis by Combi et al. (2021). Post-perihelion, our results are consistent with independent measurements obtained via infrared spectroscopy (IRTF/Keck; Faggi et al., 2021), ground-based narrow-band photometry (HCT; Aravind et al., 2025), and radio observations (NRT; Drozdovskaya et al., 2023). Similar to 3I/ATLAS, C/2020 F3 exhibited large heliocentric radial velocities ( km s-1). Consequently, accurate correction for the Swings effect, accounting for the Doppler shift of the solar Lyman- line, was critical for determining precise -factors. The discrepancy with the lower ground-based values (Aravind et al., 2025) is consistent with the aperture effects discussed in Section III.2, alongside differences in model scale lengths and computation methods identified by those authors.

### v2

@@ Abstract
3I/ATLAS is an interstellar object whose activity provides critical insights into its composition and origin. However, due to its orbital geometry, the object is too close to the Sun near perihelion to be observed from the ground, and space-based measurements are therefore required. Here we characterize the water production rate of 3I/ATLAS using SOHO/SWAN Lyman- observations from 2025 November to December (heliocentric distances 1.4 to 2.2 au) with 3D Monte Carlo modeling. We report a peak post-perihelion water production rate of molecules s-1, corresponding to a minimum active fraction of 30% (assuming a maximum nucleus radius of 2.8 km). Comparison of our post-perihelion measurements with published pre-perihelion results reveals a heliocentric asymmetry, with an scaling for the inbound rise, followed by a shallower scaling during the outbound decline, where is heliocentric distance. The post-perihelion behavior indicates that the water production of 3I/ATLAS was driven primarily by the varying solar insolation acting on a stable active area. Combined with other evidence, including comparison with the hyperactive comet 103P/Hartley 2, our findings suggest that its water production is likely dominated by a distributed source of icy grains. Furthermore, it displayed remarkable stability in the activity with no signs of outbursts or rapid depletion of water production.
@@ II.2 Monte Carlo Particle Trajectory Model (MCPTM)
The water production rate was determined by scaling the observed Lyman- integrated flux to the MCPTM predictions. Using the model described by Combi and Smyth (1988a), the relationship is expressed as:
where is the integrated flux within the photometric aperture, is the SOHO-comet distance, and is the calibration factor (Section II). is the input production rate used in the Monte Carlo simulation, and represents the modeled number of hydrogen atoms within the aperture, calculated based on the adopted velocity field () and lifetimes ().
@@ III.1 Post-perihelion phase
We derived water production rates from SOHO/SWAN observations using the MCPTM described in Section II.2. During the post-perihelion phase, the water production rate reached a peak of molecules s-1 (Table 1) and declined steadily with a heliocentric index of , as shown in Figure 3.
@@ III.2 Perihelion Asymmetry
Comparison between our post-perihelion measurements and the published pre-perihelion data reveals a significant perihelion asymmetry in water production. The activity profile is characterized by a steep pre-perihelion rise followed by a shallower post-perihelion decline (Figure 3). This asymmetry is also evident in the heliocentric light curve obtained from the COBS database111https://cobs.si (Zakrajsek and Mikuz, 2018).
@@ Appendix A Validation of the Water Production Rate Retrieval Methodology
To validate our data reduction and modeling pipeline, we benchmarked our results against three well-characterized comets: 41P/Tuttle-Giacobini-Kresák (2017 apparition), 46P/Wirtanen (2018 apparition), and C/2020 F3 (NEOWISE). These targets span nearly three orders of magnitude in water production rate ( to molecules s-1), allowing for a rigorous assessment of our analysis across a wide dynamic range. As illustrated in Figures 5, 6, and 7, our derived water production rates show excellent agreement with independent space- and ground-based measurements, confirming the robustness of our methodology.
41P/Tuttle-Giacobini-Kresák: As a typical Jupiter-family comet, 41P serves as a benchmark for the sensitivity of our analysis in the low-activity regime ( molecules s-1). As shown in Figure 5, our derived rates show excellent agreement with both the space-based Swift/UVOT results (Bodewits et al., 2018) and ground-based TRAPPIST photometry (Moulane et al., 2018). This concordance between the SWAN Lyman- (atomic hydrogen) and TRAPPIST OH data validates the photochemical scale lengths and calibration parameters adopted for low-production targets. 46P/Wirtanen:
@@
C/2020 F3 (NEOWISE): For the highly active Oort Cloud comet C/2020 F3 (NEOWISE), as shown in Figure 7, our derived water production rates closely track the heliocentric evolution reported in the standard SWAN analysis by Combi et al. (2021). Post-perihelion, our results are consistent with independent measurements obtained via infrared spectroscopy (IRTF/Keck; Faggi et al., 2021), optical spectroscopy (Himalayan Chandra Telescope (HCT); Aravind et al., 2025), and radio observations (Nançay Radio Telescope (NRT); Drozdovskaya et al., 2023). Similar to 3I/ATLAS, C/2020 F3 exhibited large heliocentric radial velocities ( km s-1). Consequently, accurate correction for the Swings effect, accounting for the Doppler shift of the solar Lyman- line, was critical for determining precise -factors. The discrepancy with the lower ground-based values (Aravind et al., 2025) is consistent with the aperture effects discussed in Section III.2, alongside differences in model scale lengths and computation methods identified by those authors.